OpenStudy (anonymous):

show that k=1 sigma k tends to infinity ((1/k)-(1/k+2)) =3/2

OpenStudy (anonymous):

So, \[\sum_{k=1}^{\infty} \frac{1}{k}-\frac{1}{k+2}= \left(\frac{1}{1}-\frac{1}{3}\ \right)+\left(\frac{1}{2}-\frac{1}{4}\ \right)+\left(\frac{1}{3}-\frac{1}{5}\ \right)+\left(\frac{1}{4}-\frac{1}{6}\ \right).... \] So as you can see after some time, the third set of terms in parenthesis starts to cancel out the negative terms. As this goes on these negative terms all fade out with all the positive terms except the first two, thus, \[1+\frac{1}{2}=\frac{3}{2}\]

OpenStudy (anonymous):

heloo any one see my qus.. plz help me.

OpenStudy (anonymous):

bro can u explain now what u did?

OpenStudy (anonymous):

Ask me questions.

OpenStudy (anonymous):

whole term whrer going.?

OpenStudy (anonymous):

Term by term is going to 0. As you can see the fractions are getting smaller.

OpenStudy (anonymous):

thisi is telescoping series..?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

how can calculte the sum of this series..?>??

OpenStudy (anonymous):

Simple, but subtle. First, look at it as sets of numbers. So, look at them in groups. The first parenthesis is the first group, then the second parenthesis is the second group, and third pare...you get the point. Scroll up and look at them then write me back when you have.

OpenStudy (anonymous):

any formula calculate the telescoping series sum?

OpenStudy (anonymous):

Yes, add all the terms up until one less than the telescoping term.

OpenStudy (anonymous):

There's a formula, but it looks ugly.

OpenStudy (anonymous):

in my qustion how to use formula ?

OpenStudy (anonymous):

What is the telescoping term?

OpenStudy (anonymous):

i dont know..:(

OpenStudy (anonymous):

Back to square one. " Simple, but subtle. First, look at it as sets of numbers. So, look at them in groups. The first parenthesis is the first group, then the second parenthesis is the second group, and third pare...you get the point. Scroll up and look at them then write me back when you have. "

OpenStudy (anonymous):

formula.?

OpenStudy (anonymous):

Here you go: \[ \sum_{k=1}^{\infty} \frac{1}{n(n+k)}=\sum_{k=1}^{\infty} \frac{1}{n}-\frac{1}{n+k}=\frac{H_k}{k}\] Where \[H_k\] is the Kth harmonic number.

OpenStudy (anonymous):

what is 1/n waht is n.?

OpenStudy (anonymous):

\[H_k=\sum_{t=1}^k \frac{1}{t}\]

OpenStudy (anonymous):

For the first term, or group of parenthesis, its n=1 for the the second one is 2 for the third is 3 for the foutrh is.....you get the point, its all the integers and as you go on looking through the set of parenthesis you find bigger values for n.

OpenStudy (anonymous):

Hk/k Hk=1/t and what is K ?

OpenStudy (anonymous):

It defines the telescoping terms within the series....unless you want to start from square one, none of these variables and constants will make sense, even if you have the formula.

OpenStudy (anonymous):

so bro in my qus what is Hk and what is K?

OpenStudy (anonymous):

K=2 and Hk is 1+1/2=3/2

OpenStudy (anonymous):

how..?

OpenStudy (anonymous):

Magic....Seriously, if you don't start form square one it wont make sense anyway.

OpenStudy (anonymous):

plz bro explain now..:(

OpenStudy (anonymous):

I got to go, sorry.